# Radial Fast Entangling Gates Under Micromotion in Trapped-Ion Quantum Computers

### v1

@@ Appendix A Derivation of fast gate condition equations in the presence of micromotion
A.1 Dimensionless motional mode quadratures
@@
$\hat{X}_{\alpha}(t)\equiv\hat{x}_{\alpha}(t)/l_{0,\alpha}$ $=\frac{\hat{a}_{\alpha}u^{*}_{\alpha}(t)+\hat{a}^{\dagger}_{\alpha}u_{\alpha}(t)}{\sqrt{2}}\,$, (A1a)
$\hat{Y}_{\alpha}(t)\equiv l_{0,\alpha}\hat{p}_{\alpha}(t)/\hbar$ $=\frac{\hat{a}_{\alpha}\dot{u}^{*}_{\alpha}(t)+\hat{a}^{\dagger}_{\alpha}\dot{u}_{\alpha}(t)}{\sqrt{2}\omega_{\alpha}}\,.$ (A1b)
@@
$\hat{H}_{\rm mot}(t)=\sum_{\alpha}\frac{\hbar\omega_{\alpha}}{2}\left(\hat{Y}_{\alpha}^{2}+\frac{\lambda_{\alpha}(t)}{\omega_{\alpha}^{2}}\hat{X}^{2}\right)\,$, (A2)
@@
$\dot{\hat{X}}_{\alpha}$ $=\omega_{\alpha}\hat{Y}_{\alpha}\,$, (A3a)
$\dot{\hat{Y}}_{\alpha}$ $=-\frac{\lambda_{\alpha}(t)}{\omega_{\alpha}}\hat{X}_{\alpha}\,.$ (A3b)
Combining these two equations gives the (quantized) Mathieu-Hill equation, $\ddot{\hat{X}}_{\alpha}(t)+\lambda_{\alpha}(t)\hat{X}_{\alpha}(t)=0$, the solution of which is given by Eq. (A1) in terms of the Mathieu-Hill mode function Eq. (8).
A.2 Semiclassical motion due to state-dependent kicks
@@
$\hat{X}_{\alpha}(t)\rightarrow X_{\alpha}(t)$ $\equiv\langle\xi_{\alpha}|\hat{X}_{\alpha}|\xi_{\alpha}\rangle=\frac{\xi_{\alpha}u^{*}_{\alpha}(t)+\xi_{\alpha}^{*}u_{\alpha}(t)}{\sqrt{2}}\,$, (A4)
@@
$X_{\alpha}(t_{0})$ $=\frac{\xi_{\alpha}+\xi_{\alpha}^{*}}{\sqrt{2}}u_{\alpha,C}(t_{0})+i\frac{\xi_{\alpha}^{*}-\xi_{\alpha}}{\sqrt{2}}u_{\alpha,S}(t_{0})\,$, (A5a)
$\omega_{\alpha}Y_{\alpha}(t_{0})$ $=\frac{\xi_{\alpha}+\xi_{\alpha}^{*}}{\sqrt{2}}\dot{u}_{\alpha,C}(t_{0})+i\frac{\xi_{\alpha}^{*}-\xi_{\alpha}}{\sqrt{2}}\dot{u}_{\alpha,S}(t_{0})\,.$ (A5b)
@@
$\xi_{\alpha}=\frac{1}{\sqrt{2}}\frac{\left(\omega_{\alpha}Y_{\alpha}(t_{0})u_{\alpha,S}(t_{0})-X_{\alpha}(t_{0})\dot{u}_{\alpha,S}(t_{0})\right)+i\left(X_{\alpha}(t_{0})\dot{u}_{\alpha,C}(t_{0})-\omega_{\alpha}(t_{0})Y_{\alpha}(t_{0})u_{\alpha,C}(t_{0})\right)}{u_{\alpha,S}(t_{0})\dot{u}_{\alpha,C}(t_{0})-u_{\alpha,C}(t_{0})\dot{u}_{\alpha,S}(t_{0})}\,.$ (A6)
@@
$X_{\alpha}(t)=\sum_{j=1}^{\mathcal{N}}\Theta(t_{j}-t)z_{j}\sqrt{2}\hat{\eta}_{\alpha}^{(A,B)}\frac{\left(u_{\alpha,S}(t-t_{j})u_{\alpha,S}(t_{j})-u_{\alpha,C}(t-t_{j})u_{\alpha,C}(t_{j})\right)}{\rho_{\alpha}(t_{j})}$ (A7)
where we have defined $\rho_{\alpha}(t_{j})=[u_{c}(t_{j})u_{s}^{\prime}(t_{j})-u_{s}(t_{j})u_{c}^{\prime}(t_{j})]/\omega_{\alpha}$.
@@
$f_{\alpha,S}(t)$ $\equiv\sum_{n}C_{n}\sin(n\Omega_{\rm RF}t+n\phi)\,$, (A8)
$f_{\alpha,C}(t)$ $\equiv\sum_{n}C_{n}\cos(n\Omega_{\rm RF}t+n\phi)$ (A9)
@@
$u_{\alpha,S}(t)$ $=\sin(\omega_{\alpha}t)f_{\alpha,C}(t)+\cos(\omega_{\alpha}t)f_{\alpha,S}(t)\,$, (A10)
$u_{\alpha,C}(t)$ $=\cos(\omega_{\alpha}t)f_{\alpha,C}(t)-\sin(\omega_{\alpha}t)f_{\alpha,S}(t)\,.$ (A11)
@@
$X_{\alpha}(t)=\sum_{j=1}^{\mathcal{N}}\Theta(t_{j}-t)\frac{z_{j}\sqrt{2}\hat{\eta}_{\alpha}^{(A,B)}}{\rho_{j}}\left[\cos(\omega_{\alpha}[t-t_{j}])\mu^{(s)}_{\alpha}(t,t_{j})+\sin(\omega_{\alpha}[t-t_{j}])\mu^{(c)}_{\alpha}(t,t_{j})\right]\,$, (A12)
in terms of the symmetric (cosine-like) and anti-symmetric (sine-like) tensors $\mu^{(c)}(t,t^{\prime})$ and $\mu^{(s)}(t,t^{\prime})$ defined in Equation (20) of the main text. The analytic solution for the momentum quadrature of each mode is simply given in terms of the time derivative of the above solution by Eq. (A3b), i.e. $Y_{\alpha}(t)=\dot{X}_{\alpha}(t)/\omega_{\alpha}$. To obtain Eq. (18) of the main text, we evaluate the mode quadratures at the end of the gate $t=t_{\rm g}$ and separate out a factor of $(b_{\alpha}^{(A)}\hat{\sigma}_{z}^{(A)}+b_{\alpha}^{(B)}\hat{\sigma}_{z}^{(B)})$ from $\hat{\eta}_{\alpha}^{(A,B)}$, i.e. defining $\Delta X_{\alpha}$ such that
$X_{\alpha}(t_{g})=\Delta X_{\alpha}(b_{\alpha}^{(A)}\hat{\sigma}_{z}^{(A)}+b_{\alpha}^{(B)}\hat{\sigma}_{z}^{(B)})\,$, (A13)
@@
A.3 State-dependent phase accumulation during gate operation
@@
Treating the position quadratures of each motional mode semiclassically – i.e. replacing $\hat{X}_{\alpha}$ with its expectation value $X_{\alpha}$ – the Lagrangian for the two-ion system is given by $\mathcal{L}=\sum_{\alpha={\rm CM,BR}}\mathcal{L}_{\alpha}$, where
$\mathcal{L}_{\alpha}(X_{\alpha},\dot{X}_{\alpha},t)=\frac{m}{2}\left(\dot{X}_{\alpha}^{2}-\lambda_{\alpha}(t)X^{2}\right)l_{0,\alpha}^{2}=\frac{\hbar}{2\omega_{\alpha}}\left(\dot{X}_{\alpha}^{2}-\lambda_{\alpha}(t)X^{2}\right)$ (A14)
@@
$\phi_{\alpha}(t_{i},t_{f})\equiv\int_{t_{i}}^{t_{f}}dt\;\mathcal{L}_{\alpha}(X_{\alpha},\dot{X}_{\alpha},t)\,.$ (A15)
@@
$\phi_{\alpha}(t_{i},t_{f})=\frac{1}{2\omega_{\alpha}}\left[X_{\alpha}(t)\dot{X}_{\alpha}(t)\right]_{t_{i}}^{t_{f}}\,.$ (A16)
@@
$\phi_{{\rm G},\alpha}=\sum_{j=1}^{\mathcal{N}}\phi_{\alpha}(t_{j}^{+},t_{j+1}^{-})\,$, (A17)
@@
$\phi_{{\rm free},\alpha}=\phi_{\alpha}(t_{1}^{+},t_{\mathcal{N}+1}^{-})\,.$ (A18)
@@
$\phi_{\alpha}=\phi_{{\rm G},\alpha}-\phi_{{\rm free},\alpha}=\frac{1}{2}\sum_{j=1}^{\mathcal{N}}X_{\alpha}(t_{j})\delta Y_{\alpha,j}\,.$ (A19)
@@
$\delta Y_{\alpha,j}\equiv\frac{\dot{X}_{\alpha}(t_{j}^{+})-\dot{X}_{\alpha}(t_{j}^{-})}{\omega_{\alpha}}=2^{3/2}z_{j}\eta_{\alpha}(b_{\alpha}^{(A)}\hat{\sigma}_{z}^{(A)}+b_{\alpha}^{(B)}\hat{\sigma}_{z}^{(B)})\,.$ (A20)
Then, substituting Equation (A12) gives the result:
$\phi_{\alpha}=8\eta_{\alpha}^{2}\hat{\sigma}_{z}^{(A)}\otimes\hat{\sigma}_{z}^{(B)}b_{\alpha}^{(A)}b_{\alpha}^{(B)}\sum_{n=1}^{\mathcal{N}}\sum_{m=2}^{n-1}\frac{z_{n}z_{m}}{\rho(t_{m})}\left[\sin(\omega_{\alpha}\Delta t_{nm})\mu_{\alpha}^{(c)}(t_{n},t_{m})+\cos(\omega_{\alpha}\Delta t_{nm})\mu_{\alpha}^{(s)}(t_{n},t_{m})\right]$ (A21)
@@
$\Theta\,\hat{\sigma}_{z}^{(A)}\otimes\hat{\sigma}_{z}^{(B)}\equiv\sum_{\alpha}\phi_{\alpha}\,$, (A22)
@@
Appendix B Multi-Modal Mathieu Equation
In this Appendix, we derive the Mathieu equation that governs the classical motion of $N=2$ trapped ions along the radial axis of a linear Paul trap.
We begin with the energetic potential for $N>1$ ions in a linear Paul trap,
[…]
Because this axis exhibits no oscillatory behavior (and, as a consequence, no micromotion), the ‘equilibrium’ positions along the $z$-axis, $\vec{\tilde{z}}(t)$, will exactly be the static equilibrium positions of $V(\vec{z})$. These static equilibrium solutions can be found via the solution to the following condition,
$\frac{\partial V(z)}{\partial z_{i}}\Biggr|_{\tilde{z}_{i}}=0\,.$ (A25)
The equilibrium solution for the radial positions of the ions lie on the RF null, i.e. $\tilde{x}_{i}=0$.
We then proceed by Taylor expanding the classical potential in terms of small transverse displacements around the static equilibria of the ions, i.e.
$V(\vec{x}(t),t)=\frac{1}{2}\sum_{i,j=0}^{N-1}\frac{\partial^{2}V}{\partial x^{(i)}\partial x^{(j)}}\Biggl|_{\vec{\tilde{x}}(t)}x^{(i)}(t)x^{(j)}(t)+\mathcal{O}\left([x^{(i)}(t)]^{3}\right)\,$, (A26)
where we have set the zero of the potential at the RF null, $x=0$. The resulting classical equations of motion can be expressed in matrix form,
$\ddot{\vec{x}}(t)=-\textbf{H}\,\vec{x}(t)\,$, (A27)
where the Hessian H has components
$[\textbf{{H}}]_{i,j}=\frac{1}{m}\frac{\partial^{2}V}{\partial x^{(i)}\partial x^{(j)}}\,.$ (A28)
Explicitly,
H $=\frac{1}{4}m\Omega_{\rm{RF}}^{2}\left(a_{x}-2q_{x}\cos(\Omega_{\rm{RF}}t)\right)\times\mathbb{1}+\frac{e^{2}}{4\pi\epsilon_{0}}\begin{bmatrix}\sum_{k\neq 0}^{N-1}\frac{-1}{\left(\tilde{z}_{0}-\tilde{z}_{k}\right)^{3}}&\dots&\frac{1}{\left(\tilde{z}_{0}-\tilde{z}_{j}\right)^{3}}&\dots&\frac{1}{\left(\tilde{z}_{0}-\tilde{z}_{N-1}\right)^{3}}\\ \vdots&\ddots&\vdots&\ddots&\vdots\\ \frac{1}{\left(\tilde{z}_{i}-\tilde{z}_{0}\right)^{3}}&\dots&\sum_{k\neq i}^{N-1}\frac{-1}{\left(\tilde{z}_{i}-\tilde{z}_{k}\right)^{3}}&\dots&\frac{1}{\left(\tilde{z}_{i}-\tilde{z}_{N-1}\right)^{3}}\\ \vdots&\ddots&\vdots&\ddots&\vdots\\ \frac{1}{\left(\tilde{z}_{N-1}-\tilde{z}_{0}\right)^{3}}&\dots&\frac{1}{\left(\tilde{z}_{N-1}-\tilde{z}_{j}\right)^{3}}&\dots&\sum_{k\neq N-1}^{N-1}\frac{-1}{\left(\tilde{z}_{N-1}-\tilde{z}_{k}\right)^{3}}\end{bmatrix}\,$.
Diagonalisation of the Hessian matrix allows us to move to a decoupled motional modes description of the dynamics in terms of its time-dependent eigenvalues $\Lambda_{\alpha}(t)$ and (normalized) time-independent eigenvectors $\vec{b}_{\alpha}$, i.e.
$\ddot{x}_{\alpha}(t)=-\Lambda_{\alpha}(t)x_{\alpha}^{2}\,$, (A29)
where $x_{\alpha}(t)=\sum_{i}b_{\alpha}^{(i)}x_{i}(t)$ is the position of the $\alpha$-th motional mode. From the above expression for the Hessian, it is clear that the eigenvalues of the first term are $\frac{1}{4}m\Omega_{\rm{RF}}^{2}\left(a_{x}-2q_{x}\cos(\Omega_{\rm{RF}}t)\right)$. The eigenvalues of the second matrix can be computed analytically for $N=2$ ions using the symmetry about the center of the trap $\tilde{z}_{1}-\tilde{z}_{0}=d$, which gives the result
$\Lambda_{\rm CM}(t)$ $=\frac{1}{4}m\Omega_{\rm{RF}}^{2}\left(a_{x}-2q_{x}\cos(\Omega_{\rm{RF}}t)\right)\,$, (A30a)
$\Lambda_{\rm BR}(t)$ $=\frac{1}{4}m\Omega_{\rm{RF}}^{2}\left(a_{x}-2q_{x}\cos(\Omega_{\rm{RF}}t)\right)+\frac{e^{2}}{\epsilon_{0}\pi d^{3}}\,.$ (A30b)
The correction arising from the Coulomb interaction between ions can be absorbed into the definition of $a_{x}$, which then becomes the mode-dependent quantity $a_{\alpha}=a_{x}+4\Gamma_{\alpha}/(m\Omega_{\rm RF}^{2})$, with $\Gamma_{\rm CM}=0$ and $\Gamma_{\rm BR}=e^{2}/(4\pi\epsilon_{0}d^{3})$.
Appendix C Micromotion Optimisation Parameters
In this Appendix, we outline how the trap parameters were defined to enable comparison between different micromotion environments during gate searches.
For gate searches presented in the main text, we consider systems of two ions confined in a linear Paul trap, where the SDKs are implemented along the radial $x$-dimension of the trap using two-photon Raman transitions. There are only two secular motional modes that are excited by the SDKs: the radial common-motional mode, $\omega_{\rm{CM}}$, and the radial breathing (stretch) mode $\omega_{\rm{BR}}$. The corresponding mode frequencies are given by,
$\omega_{\rm{CM}}=\frac{\beta_{\rm CM}\Omega_{\rm{RF}}}{2},\quad\omega_{\rm{BR}}=\frac{\beta_{\rm BR}\Omega_{\rm{RF}}}{2}\,$, (A31)
where $\beta_{\rm CM}$ and $\beta_{\rm BR}$ are the characteristic exponents of the respective Mathieu equations that govern the evolution the center-of-mass and breathing modes, respectively.
To sensibly compare different micromotion amplitudes ($q_{x}$ values) in this study, we fix the secular common-motional frequency $\omega_{\rm{CM}}$ and keep it constant across all micromotion amplitudes. We also fix the relative splitting between the modes $\chi$, which sets $\omega_{\rm{BR}}$ via,
$\omega_{\rm{BR}}=(1+\chi)\omega_{\rm{CM}}\,.$ (A32)
That is, we consider different trapping configurations that have the same secular limit for different $q_{x}$ values in the range $|q_{x}|<1$.

### v2

@@ Appendix B Derivation of fast gate condition equations in the presence of micromotion
B.1 Dimensionless motional mode quadratures
@@
$\hat{X}_{\alpha}(t)\equiv\hat{x}_{\alpha}(t)/l_{0,\alpha}$ $=\frac{\hat{a}_{\alpha}u^{*}_{\alpha}(t)+\hat{a}^{\dagger}_{\alpha}u_{\alpha}(t)}{\sqrt{2}}\,$, (26a)
$\hat{Y}_{\alpha}(t)\equiv l_{0,\alpha}\hat{p}_{\alpha}(t)/\hbar$ $=\frac{\hat{a}_{\alpha}\dot{u}^{*}_{\alpha}(t)+\hat{a}^{\dagger}_{\alpha}\dot{u}_{\alpha}(t)}{\sqrt{2}\omega_{\alpha}}\,.$ (26b)
@@
$\hat{H}_{\rm mot}(t)=\sum_{\alpha}\frac{\hbar\omega_{\alpha}}{2}\left(\hat{Y}_{\alpha}^{2}+\frac{\lambda_{\alpha}(t)}{\omega_{\alpha}^{2}}\hat{X}^{2}\right)\,$, (27)
@@
$\dot{\hat{X}}_{\alpha}$ $=\omega_{\alpha}\hat{Y}_{\alpha}\,$, (28a)
$\dot{\hat{Y}}_{\alpha}$ $=-\frac{\lambda_{\alpha}(t)}{\omega_{\alpha}}\hat{X}_{\alpha}\,.$ (28b)
Combining these two equations gives the (quantized) Mathieu-Hill equation, $\ddot{\hat{X}}_{\alpha}(t)+\lambda_{\alpha}(t)\hat{X}_{\alpha}(t)=0$, the solution of which is given by Eq. (26) in terms of the Mathieu-Hill mode function Eq. (22).
B.2 Semiclassical motion due to state-dependent kicks
@@
$\hat{X}_{\alpha}(t)\rightarrow X_{\alpha}(t)$ $\equiv\langle\xi_{\alpha}|\hat{X}_{\alpha}|\xi_{\alpha}\rangle=\frac{\xi_{\alpha}u^{*}_{\alpha}(t)+\xi_{\alpha}^{*}u_{\alpha}(t)}{\sqrt{2}}\,$, (29)
@@
$X_{\alpha}(t_{0})$ $=\frac{\xi_{\alpha}+\xi_{\alpha}^{*}}{\sqrt{2}}u_{\alpha,C}(t_{0})+i\frac{\xi_{\alpha}^{*}-\xi_{\alpha}}{\sqrt{2}}u_{\alpha,S}(t_{0})\,$, (30a)
$\omega_{\alpha}Y_{\alpha}(t_{0})$ $=\frac{\xi_{\alpha}+\xi_{\alpha}^{*}}{\sqrt{2}}\dot{u}_{\alpha,C}(t_{0})+i\frac{\xi_{\alpha}^{*}-\xi_{\alpha}}{\sqrt{2}}\dot{u}_{\alpha,S}(t_{0})\,.$ (30b)
@@
$\xi_{\alpha}=\frac{1}{\sqrt{2}}\frac{\left(\omega_{\alpha}Y_{\alpha}(t_{0})u_{\alpha,S}(t_{0})-X_{\alpha}(t_{0})\dot{u}_{\alpha,S}(t_{0})\right)+i\left(X_{\alpha}(t_{0})\dot{u}_{\alpha,C}(t_{0})-\omega_{\alpha}(t_{0})Y_{\alpha}(t_{0})u_{\alpha,C}(t_{0})\right)}{u_{\alpha,S}(t_{0})\dot{u}_{\alpha,C}(t_{0})-u_{\alpha,C}(t_{0})\dot{u}_{\alpha,S}(t_{0})}\,.$ (31)
@@
$X_{\alpha}(t)=\sum_{j=1}^{\mathcal{N}}$ $\frac{H(t_{j}-t)z_{j}\sqrt{2}\hat{\eta}_{\alpha}^{(A,B)}}{\rho_{\alpha}(t_{j})}\big[u_{\alpha,S}(t-t_{j})u_{\alpha,S}(t_{j})$
$\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;-u_{\alpha,C}(t-t_{j})u_{\alpha,C}(t_{j})\big]$ (32)
where $H(t)$ is the Heaviside step function and we have defined $\rho_{\alpha}(t_{j})=[u_{c}(t_{j})u_{s}^{\prime}(t_{j})-u_{s}(t_{j})u_{c}^{\prime}(t_{j})]/\omega_{\alpha}$.
@@
$f_{\alpha,S}(t)$ $\equiv\sum_{n}C_{n}\sin(n\Omega_{\rm RF}t+n\phi)\,$, (33)
$f_{\alpha,C}(t)$ $\equiv\sum_{n}C_{n}\cos(n\Omega_{\rm RF}t+n\phi)$ (34)
@@
$u_{\alpha,S}(t)$ $=\sin(\omega_{\alpha}t)f_{\alpha,C}(t)+\cos(\omega_{\alpha}t)f_{\alpha,S}(t)\,$, (35)
$u_{\alpha,C}(t)$ $=\cos(\omega_{\alpha}t)f_{\alpha,C}(t)-\sin(\omega_{\alpha}t)f_{\alpha,S}(t)\,.$ (36)
@@
$X_{\alpha}(t)=$ $\sum_{j=1}^{\mathcal{N}}\frac{H(t_{j}-t)z_{j}\sqrt{2}\hat{\eta}_{\alpha}^{(A,B)}}{\rho_{j}}\big[\cos(\omega_{\alpha}[t-t_{j}])$
$\times\mu^{(s)}_{\alpha}(t,t_{j})+\sin(\omega_{\alpha}[t-t_{j}])\mu^{(c)}_{\alpha}(t,t_{j})\big]\,$, (37)
in terms of the symmetric (cosine-like) and anti-symmetric (sine-like) variables $\mu^{(c)}(t,t^{\prime})$ and $\mu^{(s)}(t,t^{\prime})$ defined in Equation (10) of the main text. The analytic solution for the momentum quadrature of each mode is simply given in terms of the time derivative of the above solution by Eq. (28b), i.e. $Y_{\alpha}(t)=\dot{X}_{\alpha}(t)/\omega_{\alpha}$. To obtain Eq. (8) of the main text, we evaluate the mode quadratures at the end of the gate $t=t_{\rm g}$ and separate out a factor of $(b_{\alpha}^{(A)}\hat{\sigma}_{z}^{(A)}+b_{\alpha}^{(B)}\hat{\sigma}_{z}^{(B)})$ from $\hat{\eta}_{\alpha}^{(A,B)}$, i.e. defining $\Delta X_{\alpha}$ such that
$X_{\alpha}(t_{g})=\Delta X_{\alpha}(b_{\alpha}^{(A)}\hat{\sigma}_{z}^{(A)}+b_{\alpha}^{(B)}\hat{\sigma}_{z}^{(B)})\,$, (38)
@@
B.3 State-dependent phase accumulation during gate operation
@@
Treating the position quadratures of each motional mode semiclassically – i.e. replacing $\hat{X}_{\alpha}$ with its expectation value $X_{\alpha}$ – the Lagrangian for the two-ion system is given by $\mathcal{L}=\sum_{\alpha={\rm CM,SR}}\mathcal{L}_{\alpha}$, where
$\mathcal{L}_{\alpha}(X_{\alpha},\dot{X}_{\alpha},t)=\frac{m}{2}\left(\dot{X}_{\alpha}^{2}-\lambda_{\alpha}(t)X^{2}\right)l_{0,\alpha}^{2}$ (39)
@@
$\phi_{\alpha}(t_{i},t_{f})\equiv\int_{t_{i}}^{t_{f}}dt\;\mathcal{L}_{\alpha}(X_{\alpha},\dot{X}_{\alpha},t)\,.$ (40)
@@
$\phi_{\alpha}(t_{i},t_{f})=\frac{1}{2\omega_{\alpha}}\left[X_{\alpha}(t)\dot{X}_{\alpha}(t)\right]_{t_{i}}^{t_{f}}\,.$ (41)
@@
$\phi_{{\rm G},\alpha}=\sum_{j=1}^{\mathcal{N}}\phi_{\alpha}(t_{j}^{+},t_{j+1}^{-})\,$, (42)
@@
$\phi_{{\rm free},\alpha}=\phi_{\alpha}(t_{1}^{+},t_{\mathcal{N}+1}^{-})\,.$ (43)
@@
$\phi_{\alpha}=\phi_{{\rm G},\alpha}-\phi_{{\rm free},\alpha}=\frac{1}{2}\sum_{j=1}^{\mathcal{N}}X_{\alpha}(t_{j})\delta Y_{\alpha,j}\,.$ (44)
@@
$\delta Y_{\alpha,j}$ $\equiv\frac{\dot{X}_{\alpha}(t_{j}^{+})-\dot{X}_{\alpha}(t_{j}^{-})}{\omega_{\alpha}}$ (45)
$=2^{3/2}z_{j}\eta_{\alpha}(b_{\alpha}^{(A)}\hat{\sigma}_{z}^{(A)}+b_{\alpha}^{(B)}\hat{\sigma}_{z}^{(B)})\,.$
Then, substituting Equation (37) gives the result:
$\phi_{\alpha}=8\eta_{\alpha}^{2}\hat{\sigma}_{z}^{(A)}\otimes\hat{\sigma}_{z}^{(B)}b_{\alpha}^{(A)}b_{\alpha}^{(B)}\sum_{n=1}^{\mathcal{N}}\sum_{m=2}^{n-1}\frac{z_{n}z_{m}}{\rho(t_{m})}\left[\sin(\omega_{\alpha}\Delta t_{nm})\mu_{\alpha}^{(c)}(t_{n},t_{m})+\cos(\omega_{\alpha}\Delta t_{nm})\mu_{\alpha}^{(s)}(t_{n},t_{m})\right]$ (46)
@@
$\Theta\,\hat{\sigma}_{z}^{(A)}\otimes\hat{\sigma}_{z}^{(B)}\equiv\sum_{\alpha}\phi_{\alpha}\,$, (47)
@@
Appendix C Choice of trapping parameters
In this Appendix, we outline how the trap parameters were determined to sensibly compare different micromotion amplitudes ($q_{x}$ values) in this study. We chose to vary the axial trapping strength to ensure that the system dynamics share the same secular limit across all $q_{x}$ values considered in this study. As we have reported gate times in units of the radial centre-of-mass frequency, $\omega_{0}\equiv\omega_{\rm CM}$, it is sufficient to fix the relative splitting between the modes $\chi=\omega_{\rm{SR}}/\omega_{\rm{CM}}-1$.